OpenStudy (anonymous):

Plz Help. Will give medal and fan. Suppose you have a wallet with $5 bills, $10 bills, and $20 bills. If the probability of picking a $20 bill is 4/15, and the probability of picking a $20 bill is 4/15, and the probability of picking a $5 bill is 1/3, what is the probability of picking a $20 bill?

OpenStudy (anonymous):

Choices are- 1/15 1/5 4/15 2/5

jimthompson5910 (jim_thompson5910):

It gives you " If the probability of picking a $20 bill is 4/15...and the probability of picking a $20 bill is 4/15" then it asks "what is the probability of picking a $20 bill?" Something seems wrong here. It seems like this is a trick question.

OpenStudy (anonymous):

I know. And I'm not good at Probability... Can you help me?

OpenStudy (anonymous):

No explained it wrong

jimthompson5910 (jim_thompson5910):

It literally tells you the probability of picking a $20

jimthompson5910 (jim_thompson5910):

which is why I'm thinking something is off here

OpenStudy (anonymous):

Suppose you have a wallet with $5 bills, $10 bills, and $20 bills. If the probability of picking a $10 bill is 4/15, and the probability of picking a $5 bill is 1/3, what is the probability of picking a $20 bill? Sorry that's the right question

jimthompson5910 (jim_thompson5910):

oh ok

jimthompson5910 (jim_thompson5910):

what do you get when you add 4/15 and 1/3 ?

OpenStudy (anonymous):

3/5

jimthompson5910 (jim_thompson5910):

so the probability of picking a $5 or a $10 is 3/5 let x = probability of picking a $20 ------------------------------------------------------- (probability of picking a $5 or $10) + (probability of picking a $20) = 100% 3/5 + x = 1 solve for x to get x = ???

OpenStudy (anonymous):

2/5?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

So the answer is 2/5?

OpenStudy (anonymous):

Thank you

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

np